how do you solve: a=1/2b+1
a-2b=-2

Answer by deepikabansal(5)

a=1/2b+1
a-2b=-2
you have two equations and two variables a and b
so from first equation we have
a=1/2b+1
=>take lcm as 2b on left hand side
=>a=(1+2b)/2b
cross multiply
=>a*2b=1+2b(consider it equation 1)
from second equation we have
a-2b=-2
=> -2b=-2-a
=>take minus sign common on both sides
2b=2+a
put this value of 2b in equqtion 1
we have
a*(2+a)=1+2+a
=>2a+a^2=3+a
=>a^2=3+a-2a
=>a^2=3-a
=>a^2+a-3=0
consider it as quadratic equation
we have two values of a
a1=(-1+(1-4*(-3)))/2
a1=(-1+(13^1/2))/2
and a2=(-1-(13^1/2))/2


by deepika bansal
mail at bdeepika11@rediffmail.com